A 25.0-kg boy stands 2.00 m from the center of a frictionless playground merry-goround, which has a moment of inertia of 200. . The boy begins to run in a circular path with a speed of 0.600 ⁄ relative to the ground.


1
Expert's answer
2021-03-31T16:14:51-0400

Solution.

"m_b=25.0kg;"

"R=2.00m;"

"I=200kgm^2;"

"v_b=0.6m\/s;"

"\omega-?;"

"v_{1b}-?;"

"I=\dfrac{1}{2}mR^2\implies m=\dfrac{2I}{R^2}" ;

"m=\dfrac{2\sdot 200kgm^2}{4.00m^2}=100kg;"

"m_bv_b=mv\implies v=\dfrac{m_bv_b}{m};"

"v=\dfrac{25.0kg\sdot 0.6m/s}{100kg}=0.15m/s;"

"\omega =\dfrac{v}{R};"

"\omega=\dfrac{0.15 m/s}{2m}=0.075s^{-1};"

"v_{1b}=v_b-v;"

"v_{1b}=0.6m\/s-0.15m\/s=0.45m\/s;"


Answer: "\omega=0.075s^{-1};" "v_{1b}=0.45m\/s."